OpenStudy (anonymous):

does [(-1)^2n ln(n)]/n converge, converge absolutely, or diverge?

OpenStudy (anonymous):

terms go to zero, and it alternates

OpenStudy (anonymous):

i assume again you mean the sum if it alternates, and the terms go to zero, that is enough for convergence

OpenStudy (anonymous):

sure as hell won't converge absolutely though, because without the \((-1)^n\) you get \[\sum\frac{\ln(n)}{n}\]

OpenStudy (anonymous):

here the degree of the denominator is only 1, so it is larger than \[\sum\frac{1}{n}\]which does not converge

OpenStudy (anonymous):

\((-1)^{2n}=1\) for all \(n\), unless that's a mistake in your series. If it's not, you have \(\dfrac{\ln n}{n}\), where \(\ln n<n\) for \(n>1\), and thus a monotonically decreasing sequence of positive terms. Apply the integral test. \[\sum_{n=1}^\infty \frac{\ln n}{n}~\text{converges/diverges if}~\int_1^\infty \frac{\ln x}{x}~dx~\text{converges/diverges, respectively.}\]

OpenStudy (anonymous):

oh lord i am blind

OpenStudy (anonymous):

on the other hand, in my defence, who writes \((-1)^{2n}\)??

OpenStudy (anonymous):

Beats me... which is why I think it might be a mistake, or a test.

OpenStudy (anonymous):

i wouldn’t use the integral test though, i would use the eyeball test

OpenStudy (anonymous):

So how would you solve it?

OpenStudy (anonymous):

If you're using the integral test, compute the integral. I'm not familiar with the eyeball test myself...

OpenStudy (anonymous):

\[\frac{\ln(n)}{n}>\frac{1}{n}\]

OpenStudy (anonymous):

^^^There you go. By the way, even if it is meant to say (-1)^n, WA says it converges... http://www.wolframalpha.com/input/?i=Sum%5B%28-1%29%5En*Log%5Bn%5D%2Fn%2C%7Bn%2C1%2CInfinity%7D%5D

OpenStudy (anonymous):

eyeball test: denominator is a polynomial of degree only 1, so it does not converge